OpenStudy (angel_kitty12):

Help ASAP I really appreciate it. I such a probability and need to learn this. A bag of marbles contains 12 red marbles, 8 blue marbles, and 5 green marbles. If three marbles are pulled out, find each of the following probabilities. In each we specify either replacement (the marbles go back into the bag after each pull) or no replacement. A. Find the probability of pulling out 3 marbles of the same color without replacement. (More complex) D. Find the probability of pulling out two blue marbles and one green marble in any order with replacement. Be careful as there are multiple ways this

OpenStudy (jack1):

so @Angel_Kitty12 will work through this with you: Q1. how many marbles in the bag total at the start? Q2. probability the first marble pulled out is a red? Q3. probability the first marble pulled out is a blue? Q4. probability the first marble pulled out is a green?

OpenStudy (angel_kitty12):

25 marbles in total. There are 12 red and the probability of them being replaced is 0.110592 There are 8 blue so that total will be 0.032768 And there are 5 green so 0.008 I think unsure here

OpenStudy (angel_kitty12):

First marble oh wait

OpenStudy (angel_kitty12):

If the first marble pulled out is red, would the fraction form be 12/25 still?

OpenStudy (angel_kitty12):

And with the blue and green marble its 8/25 and 5/25. Right? Then do I multiply them together or.... I'm very confused. So sorry

OpenStudy (jack1):

25 total = correct A: same colour (without replacement) RED = 12/25 x 11/24 x 10/23 = ? GREEN = 5/25 x 4/24 x 3/23 = ? BLUE = 8/25 x 7/24 x 6/23 = ?

OpenStudy (angel_kitty12):

Oh I read the question wrong! I thought it said with replacement! No wonder I wasn't understanding this

OpenStudy (jack1):

all good... so P(RED) = ?

OpenStudy (angel_kitty12):

So I multiply them together and add all three up to get approximately 0.16 or 0.155556174? Red= approximately 0.096 Blue= 0.021504 Green= 0.0384 Right?

OpenStudy (jack1):

i got .0244 for blue?

OpenStudy (mathmate):

@angel_kitty12 In probability it is preferable to work with fractions because they are exact, while decimals need to be rounded off, so inexact.

OpenStudy (angel_kitty12):

Or do I multiply all three after finding the sum

OpenStudy (angel_kitty12):

Okay so to work with fractions instead I got 336/13800

OpenStudy (angel_kitty12):

For blue

OpenStudy (angel_kitty12):

Which is 0.02434

OpenStudy (angel_kitty12):

I multiplied wrong your right

OpenStudy (jack1):

.0957 = red = 11/115 .0044 = blue = 1/230 .0244 = green = 14/575

OpenStudy (jack1):

cools :) all good then

OpenStudy (angel_kitty12):

Okay 0.096 or 1320/13800 for red 0.0244 or 336/13800 for blue and 0.0043 or 60/13800 for green

OpenStudy (angel_kitty12):

But I rounded wrong okay I got it now

OpenStudy (jack1):

also @Angel_Kitty12 be careful what you say around mathmale, the guy tends to abuse the ban powers he has as a moderator, so dont break the code of conduct if u see him online or in ur post

OpenStudy (angel_kitty12):

Now what's the next step? Do I add them all or multiply them all?

OpenStudy (jack1):

add them

OpenStudy (angel_kitty12):

Oh okay. That's odd

OpenStudy (jack1):

p(3 of 1 colour) = P (3red) + P(3 green) + P(3 blue)

OpenStudy (angel_kitty12):

Okay I got 0.1245

OpenStudy (jack1):

= 143 / 1150 = .1244... so yeah, perfect!

OpenStudy (angel_kitty12):

Okay cool! Now for part b though.. Will I do something similar to this?

OpenStudy (jack1):

i only see parts A and D posted...?

OpenStudy (angel_kitty12):

D I mean. That's supposed to be B

OpenStudy (angel_kitty12):

Autocorrect

OpenStudy (jack1):

ah, cool so, what are the combinations that give us 2 blue and 1 green?

OpenStudy (jack1):

as i see it: G B B B G B B B G yeah?

OpenStudy (angel_kitty12):

Okay 2 blue and one green. I know the factions are 8/25 and for green its 5/25 If it's replaceable, will it be P(8/25•8/25)+ P(5/25)?

OpenStudy (angel_kitty12):

Oh I thought something else never mind I see what your doing

OpenStudy (jack1):

no i think ur right, i didnt notice that this one was replaced each time

OpenStudy (jack1):

P(8/25) * P(8/25) * P(5/25) + P(8/25) * P(5/25) *P(8/25) + P(5/25) * P(8/25) * P(8/25) ... is that right?

OpenStudy (angel_kitty12):

Wait I'll have to add the equations three times?

OpenStudy (jack1):

well... yeah... i think so? otherwise you're just working out the probability for the ones taken out in that order only i think?

OpenStudy (angel_kitty12):

Oh okay I get why

OpenStudy (angel_kitty12):

Well it seems right I mean it's reasonable

OpenStudy (angel_kitty12):

So when I work it out, I get 0.9072 correct?